OpenStudy (anonymous):

Consider the line -6x+8y=9 . Find the equation of the line that is perpendicular to this line and passes through the point -3,-2 Find the equation of the line that is parallel to this line and passes through the point -3,-2

OpenStudy (anonymous):

Convert the line to slope intercept form

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

-6x+8y=9 8y=6x+9

OpenStudy (anonymous):

\[y=\frac{ 3 }{ 4 }x+\frac{ 9 }{ 8 }\]

OpenStudy (anonymous):

The slope would be 3/4

OpenStudy (anonymous):

okay i got -3/4 ?

OpenStudy (anonymous):

What would be the slope of a line perpendicular to this line?

OpenStudy (anonymous):

It's not negative..

OpenStudy (anonymous):

I see where i got it wrong

OpenStudy (anonymous):

ok now a line that is perpendicular has the negative reciprocal of this slope which is -4/3

OpenStudy (anonymous):

where did you get -4/3 ???

OpenStudy (anonymous):

ok let's go a little slower..

OpenStudy (anonymous):

you flipped the equation ? but how does it become negative ?

OpenStudy (anonymous):

perpendicular lines have the negative opposite of the other slope.. Example: What is the perpendicular slope of 2?

OpenStudy (anonymous):

-2 ?

OpenStudy (anonymous):

Close.. -1/2

OpenStudy (anonymous):

;/

OpenStudy (anonymous):

You have to flip the number and make it negative

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Do you understand a little better?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so now that we have the perpendicular slope which is -4/3.. we can plug it into the point slope form which is \[Y-Y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

\[y+2=-\frac{ 4 }{ 3 }(x+3)\]

OpenStudy (anonymous):

solve for y